James is one and a half times taller than his sister Jill. Together their heigh is 10 ft 5 ins . What are their individual heights?
Answer by stanbon(75887)

James is one and a half times taller than his sister Jill. Together their height is 10 ft 5 ins . What are their individual heights?
----------------
Let Jill be "x" inches tall:
EQUATION:
x + (3/2)x = 127 inches
(5/2)x = 127
x = (2/5)*127
x = 50.8 inches (Jill's height)
===================
Cheers,
Stan H.
=============
Cheers,
Stan H.